OpenStudy (anonymous):

Help!! Find the amplitude y = -5 sin(1/4)x.

OpenStudy (zzr0ck3r):

you mean \(-5\sin(\frac{1}{4}x)\)?

OpenStudy (zzr0ck3r):

for \(a\sin\)(x) \(a\) is the amplitude

OpenStudy (anonymous):

well it looks like -5sin1/4x

OpenStudy (zzr0ck3r):

but the x is on the inside

OpenStudy (anonymous):

so would the - not be apart of that?

OpenStudy (zzr0ck3r):

it does not make sense to ask for amplitude otherwise...

OpenStudy (zzr0ck3r):

sorry its \(|a|\)

OpenStudy (anonymous):

ok ok, I understand because i thought the amplitude was a so i thought it was -5 but that wasn't one of my answers, thank you :)